INTERMOLECULAR INTERACTIONS BETWEEN THE SUPEROXIDE RADICAL AND HYDROGEN FLUORIDE

\underline{WAFAA M. FAWZY}, YUCHENG ZHANG, and MAHMOUD ELSAYED, Department of Chemistry, Murray State University, Murray, KY 42071.

This work concerns theoretical characterization of the nature of intermolecular interaction between the superoxide radical and hydrogen fluoride. The long-range intermolecular potential energy surface of the superoxide-HF complex was examined using the CCSD(T)/aug-cc-pVXZ, X=1, 2, 3, 4, and 5, level of theory. Our preliminary results show a minimum energy structure that corresponds to a non-linear planar geometry of Cs symmetry. In this structure, the hydrogen atom of the HF moiety is bonded to one of the oxygen atoms of the superoxide radical via a very strong hydrogen bond that is comparable to ionic bond (D_{e} = 41.4 kcal/mol). Such an ionic hydrogen bond causes elongation of the H-F bond length in the complex by about 0.1 angstroms. These results as well as those of vibrational frequency calculations will be presented.